Problem 19.47 An object with a charge of-4.5 μC and a mass of 1.5x10-2 kg experiences an upward electric force, due to a urtform eoctnc field, equal in magnitude to its weight. Part A Find the magnitude of the electric field. Express your answer using two significant figures ANSWER Part B Find the direction of the electric field. ANSWER: O upward e downward O to the left to the right
Answer #1

part A)

let the electric field is E

Now, for the magnitude of electric field

weight = electric force

m * g = e * E

1.5 *10^-2 * 9.8 = 4.5 *10^-6 * E

E = 32667 N/C

the electric field is 32667 N/C

part B)

as the electric force is upwards

as the charge is negative

the electric field is downwards

hence , the direction of electric field is Downward